Two identical point charges (q = +8.40 x 10-6 C) are fixed at opposite corners of a square whose sides have a length of 0.640 m. A test charge (q0 = -2.80 x 10-8 C), with a mass of 7.20 x 10-8 kg, is released from rest at one of the corners of the square. Determine the speed of the test charge when it reaches the center of the square.
Answer #1

here,

q = 8.4 * 10^-6 C

a = 0.64 m

q0 = - 2.8 * 10^-8 C

mass , m = 7.2 * 10^-8 kg

let the speed of test charge at the center of square be v

using Work-energy theorm

( Vcenter - Vcorner) * q0 = 0.5 * m * v^2

( (2 * K * q /sqrt(a^2 + a^2)/2) - (2 * K * q /a)) * (- 2.8 * 10^-8) = 0.5 * 7.2 * 10^-8 * v^2

( (9 * 10^9 * 8.4 * 10^-6 /sqrt(2 * 0.64^2)) - (2 * 9 * 10^9 * 8.4 * 10^-6 /0.64)) * (-2.8 * 10^-8) = 0.5 * 7.2 * 10^-8 * v^2

solving for v

v = 344.7 m/s

the speed of test charge at the center of square is 344.7 m/s